### v1

# Do LVLMs Understand Charts? Analyzing and Correcting Factual Errors in Chart Captioning

Abstract
Recent advancements in large vision-language models (LVLMs) have led to significant progress in generating natural language descriptions for visual content and thus enhancing various applications. One issue with these powerful models is that they sometimes produce texts that are factually inconsistent with the visual input. While there has been some effort to mitigate such inconsistencies in natural image captioning, the factuality of generated captions for structured document images, such as charts, has not received as much scrutiny, posing a potential threat to information reliability in critical applications. This work delves into the factuality aspect by introducing a comprehensive typology of factual errors in generated chart captions. A large-scale human annotation effort provides insight into the error patterns and frequencies in captions crafted by various chart captioning models, ultimately forming the foundation of a novel dataset, Chocolate. Our analysis reveals that even state-of-the-art models, including GPT-4V, frequently produce captions laced with factual inaccuracies. In response to this challenge, we establish the new task of Chart Caption Factual Error Correction and introduce ChartVE, a model for visual entailment that outperforms proprietary and open-source LVLMs in evaluating factual consistency. Furthermore, we propose C2TFec, an interpretable two-stage framework that excels at correcting factual errors. This work inaugurates a new domain in factual error correction for chart captions, presenting a novel evaluation mechanism, and demonstrating an effective approach to ensuring the factuality of generated chart captions. [^1] The code and data will be made publicly available: https://github.com/khuangaf/CHOCOLATE.
Do LVLMs Understand Charts?
Analyzing and Correcting Factual Errors in Chart Captioning
[…]
1 Introduction
Large vision-language models (LVLMs) have recently shown impressive capabilities in generating natural language descriptions of visual content like images, videos and charts Li et al. (2020, 2022); Wang et al. (2022); OpenAI (2023b); Google (2023a); Liu et al. (2023c); Wang et al. (2023). Chart captioning is particularly important for data analysts, business analysts, and journalists who rely on accurate chart interpretations for decision-making and reporting. However, no prior work have studied the factuality of the generated captions. Factuality, also known as the faithfulness or factual consistency between inputs and outputs, of generated captions is of critical importance when leveraging LVLMs for both practical applications and research advancements. In the domain of information dissemination, for instance, such models can be deployed to automatically generate descriptions for charts in news articles Liu et al. (2021), educational materials Fu et al. (2022), and on social media platforms Monteiro et al. (2017). Ensuring that these captions are factual directly impacts the credibility and reliability of the content provider.
To better understand the factual errors produced by chart captioning models, we first introduce a typology of factual errors particularly relevant in the chart domain. Using this scheme, we conduct a large-scale human annotation study to analyze the distributions and frequencies of various error types, such as Value Error and Label Error, across captions produced by various chart captioning models, from task-specific fine-tuned models to LVLMs (see Table 1). The annotated samples are then categorized into three splits, Lvlm (Large-vision Language Models), Llm (Large Language Models), and Ft (Fine-tuned), based on the architecture and the scale of the underlying models that generate the captions, and form a dataset which we named Chocolate. With this dataset collected, we aim to answer three main research questions. First, are state-of-the-art chart captioning models able to produce factual captions? We find the answer is no (Section 2). Out of the 1,187 captions we analyzed, only 213 of them are factual (see Table 2). Even state-of-the-art LVLM, GPT-4V OpenAI (2023b), produces a great portion of errors in its generated captions (see Figure 1). In addition, we show that different chart captioning models exhibit substantially different factual error distributions in generated captions across different chart domains.
The prevalence of factual inconsistencies observed in the generated captions by various models underscores the urgent need to address the factual accuracy of such technologies. Consequently, we introduce a task named Chart Caption Factual Error Correction (Section 3), which presents a novel challenge of rectifying factual inaccuracies in chart captions generated by LVLMs. A pertinent question that arises from this task is: how can we automatically evaluate the factual consistency between charts and captions/corrections? To tackle this question, we conceptualize the evaluation procedure as a task of visual entailment. We present the first model for chart visual entailment, named ChartVE. This model is trained by repurposing existing datasets from chart summarization and chart question answering. Our comprehensive correlation analysis shows that ChartVE performs competitively with proprietary LVLMs, GPT-4 and Bard Google (2023a), and outperforms an open-source LVLM, LLaVA-1.5 Liu et al. (2023c), despite being considerably smaller—100 times less in size. The findings emphasize the efficiency and efficacy of ChartVEn evaluating factual accuracy for the task of Chart Caption Factual Error Correction.
Figure 1: Error distribution for different models on VisText and Pew. The error rates are computed per sentence. An error rate of 0.4 indicates that 40% of the sentences in the generated captions contain such an error. Note that a single caption may contain multiple error types; hence, the maximum value for a stacked bar is greater than 1.0. We show that even the most advanced LVLM, GPT-4, generates captions with a high rate of factual error.
Now that we have set up the task, we turn to the challenge of how to effectively correct factual errors in chart captions? Here, we propose C2TFec (Section 4), an interpretable two-step framework that breaks down visual reasoning into image-to-structure rendering and text-based reasoning. C2TFec first transforms the input chart into a structured data table representation. Grounded in this extracted tabular data, the second component then identifies and fixes any factual inconsistencies in the initial caption through a transparent reasoning process. Our experiments demonstrate that this explicit decomposition enables more reliable factuality corrections compared to end-to-end approaches. The intermediate symbolic data acts as an effective bridge between charts and captions. By exploiting the complementary strengths of a specialized Chart-To-Table model and GPT-4, C2TFec significantly outperforms competitive baselines including GPT-4V, as validated through both automatic evaluation and human assessment.
In summary, our contributions are as follows:
  • We present the first empirical analysis of factual errors in captions produced by models of various scales and architectures.
  • We introduce the Chart Caption Factual Error Correction task that challenges models to correct factual errors in generated chart captions and construct the Chocolate dataset for this task by defining a novel typology of factual errors in chart captions.
  • We present ChartVE, a reference-free evaluation metric based on visual entailment that correlates better with human judges compared to LVLM-based approaches.
  • We propose C2TFec, an interpretable two-stage error correction framework that demonstrates better performance compared to all existing LVLMs.
2 Dataset Creation
To understand the capabilities of existing models in summarizing key information from charts, we conduct a large-scale analysis on six most advanced chart captioning models on the VisText Tang et al. (2023a) and Pew Kantharaj et al. (2022) datasets. To facilitate this process, we introduce an error typology, as illustrated in Section 2.1. Upon gathering human annotations, we discuss the quality of the collected data (Section 2.3) and present a detailed analysis of different captioning models (Section 2.2).
2.1 Error Typology
To understand the frequency of various types of errors made by chart captioning systems, we define a typology of errors as detailed below and demonstrate examples in Table 1.
Chart Category Example Caption
Value Error
Asians have a turnout rate of 20.4% in 1990.
Label Error
Asians have the highest turnout rates across the years.
Trend Error
From 1986-2014, the turnout rates are increasing overall.
Magnitude Error
From 1986-2014, the turnout rates are sharply decreasing overall.
Out-of-context Error
Vietnamese have the highest turnout rates among Asians.
Nonsense Error
From 1986-2014, #?sep #sep #sep #sep.
Grammatical Error
The turnout rates are decrease overall.
Table 1: Typology of errors illustrated with an example chart.
Value Error
A quantitative data value from the chart is incorrectly stated in the caption. This includes numbers representing values on axes, percentages, or other numerical data points.
Label Error
A non-numerical label, category, or text element from the chart is incorrectly referenced in the caption. This includes labels on axes, legend items, categorical variables, etc.
Trend Error
The overall direction of change over time or comparison between groups is incorrectly described in the caption, such as stating an increasing trend when the values in the chart are decreasing.
Magnitude Error
The degree or amount of difference described for a trend or comparison is factually inconsistent with the chart, such as stating a small/sharp increase/decrease when the change is more gradual based on the chart’s data.
Out-of-context Error
Concepts, variables, or any information introduced in the caption that do not exist at all in the content of the chart. The caption contains factual statements not grounded in the actual chart contents.
Nonsense Error
The caption contains incomplete sentences, disconnected phrases that do not connect logically, or sequences of words that simply do not make coherent sense.
Grammatical Error
There are grammatical mistakes in the structure or syntax of the caption. [^2] Note that we do not consider grammatical errors as factual inconsistency. We analyze grammatical errors to understand how fluent the captioning models are.
2.2 Captioning Model Analysis
We consider the various types of models. First, ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) are the most advanced task-specific models fine-tuned with in-domain data from the VisText and Pew datasets. Second, DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a) is a LLM-based pipeline approach. Finally, GPT-4V OpenAI (2023b) and Bard Google (2023a) [^3] We tested Bard before the release of Gemini Google (2023b). are the strongest LVLMs. For each model and dataset, we randomly sample 100 chart figures and generate the corresponding captions. Invalid output sequences, such as empty strings, are filtered out.
We compute the percentage of sentences with factual errors for different models and datasets, with a breakdown of the proportion of different error types. Here, we focus on the error distribution on the sentence level instead of the caption level since different models generate captions of different lengths. A sentence-level evaluation can help mitigate this discrepancy and facilitate a fairer comparison.
From Figure 1, we made the following observations. First, SOTA chart captioning models often fail to produce factual captions. Additionally, as shown in Table 2, we calculated the percentage of non-factual captions, revealing that 82.05% of captions contain at least one factual error. More importantly, even advanced models like GPT-4V and Bard, which have demonstrated proficiency in a variety of vision-language tasks, produce factually incorrect captions 81.27% of the time, as recorded in Table 5. These findings highlight the inherent difficulties of chart understanding tasks and the limitations of SOTA vision-language models.
Second, task-specific chart captioning models and LVLMs show opposite trends on the two datasets. Task-specific fine-tuned models, including ChartT5, MatCha, and UniChart, produce fewer errors on the VisText Tang et al. (2023a) dataset. Conversely, LVLMs, including GPT-4V and Bard, generate significantly fewer errors on the Pew dataset. The key distinctions on these datasets are two: (1) the vast majority of the charts in Pew datasets have corresponding numbers labeled on or next to each data point and (2) charts in VisText are associated with fewer complex structures compared to those in the Pew dataset. We hypothesize that LVLMs may be better at capturing the numbers labeled next to the data points in charts, whereas task-specific models are more capable of reading the exact value of the data points by aligning them to axes. We show an example to validate this hypothesis in Figure 6.
Third, LVLMs cannot consistently outperform task-specific fine-tuned models. Although LVLMs are trained on data with a much larger scale using many more parameters, task-specific models may perform better than LVLMs when suitable pre-training objectives and model architectures are adopted. For example, on the VisText dataset, we see that UniChart achieves a much better performance than Bard and is comparable to GPT-4V in terms of producing more factual captions. This can be attributed to UniChart’s various pre-training objectives tailored for chart understanding, enabling it to better identify the relationship between data points within the charts from VisText.
The dataset resulting from the analysis is named Chocolate (Captions Have Often Chosen Obfuscating Lies About The Evidence), where each instance consists of a chart, a generated chart caption, and error types labeled by human annotators. Drawing insights from Tang et al. (2023b) that factual errors produced by different kinds of models may be easier or more difficult to identify, we split Chocolate into three categories. First, the Lvlm split corresponds to captions produced by GPT-4V and Bard. Second, the Llm split contains captions generated by DePlot + GPT-4. Finally, the Ft split is associated with captions from ChartT5, UniChart, and MatCha. The details of each split are shown in Appendix C.
2.3 Dataset Quality
To evaluate the quality of Chocolate, we measured inter-annotator agreement by calculating Fleiss’ Kappa $\kappa$ Fleiss (1971) and the majority vote agreement percentage $p$, in line with the metrics used by Pagnoni et al. (2021). We applied these metrics across all 5,323 sentences in Chocolate. For determining factual consistency between chart sentences and their corresponding charts, we achieved a Fleiss’ Kappa of $\kappa=0.63$ and a majority vote agreement of $p=91\%$. For context, Pagnoni et al. (2021) reported a Fleiss’ Kappa of $\kappa=0.58$ and a majority agreement level of $p=91\%$. This suggests that Chocolate exhibits a quality on par with well-established benchmarks in text-based factual inconsistency detection.
# Factual # Non-factual # Total Sentence 2,561 2,762 5,323 Caption 213 974 1,187
Table 2: Statistics of the captions we analyzed. A sentence is considered factual if and only if it does not contain any factual error. A caption is considered factual if all its sentences are factual.
3 The Chart Caption Factual Error Correction Task
The dataset collected in Section 2 enables us to study the Chart Caption Factual Error Correction task. In this section, we first formally provide the definition of this task (Section 3.1) and propose an effective reference-free evaluation metric based on chart visual entailment (Section 3.2).
3.1 Task Definition
The input to our task is a chart $\mathcal{E}$ and chart caption $\mathcal{C}$ that may or may not be factually consistent with $\mathcal{E}$. The goal of chart caption factual error correction is to produce a corrected caption $\hat{\mathcal{C}}$ that fixes factual errors in $\mathcal{C}$ with the minimum amount of edits. If $\mathcal{C}$ is already faithful to $\mathcal{E}$, models should output the original caption (i.e. $\hat{\mathcal{C}}$ = $\mathcal{C}$). Following prior work on text-based factual error correction Thorne and Vlachos (2021); Huang et al. (2023b); Gao et al. (2023), corrections should be made with as few substitution, insertion, and deletion operations as possible since one can trivially achieve $0\%$ non-factual rate by deleting all words in a caption.
3.2 Reference-free Evaluation With Chart Visual Entailment
There was no established metric for evaluating the factual consistency between a chart and the corresponding chart caption. In addition, since our dataset does not contain annotated reference corrected captions [^4] We did not collect reference corrected captions due to the challenges of curating high-quality references through crowd-sourcing., text-based metrics such as SARI Xu et al. (2016) cannot be adopted. As a solution, we propose ChartVE, a reference-free evaluation metric based on chart visual entailment, as detailed in the following paragraphs.
ChartVE Overview
We formulate the task of chart caption factual inconsistency detection as a chart visual entailment task. Given a chart caption sentence $c$ and a chart $\mathcal{E}$, the task is to predict whether the relationship from $\mathcal{E}$ to $c$ as Entailment (i.e. factually consistent) or NotEntailment (factually inconsistent). The main challenge of learning a visual entailment model for this task is the lack of data. To overcome this challenge, we repurpose data from relevant tasks, such as chart QA and chart captioning, as positive samples. Then, we propose a table-guided negative data generation to produce negative samples. In the following paragraphs, we detail how each step is achieved.
Positive Data Creation
We consider datasets from two tasks that are closely related to the chart visual entailment task: chart question answering and chart captioning. The chart question answering task requires models to respond to questions regarding a specific chart, with each instance comprising a chart accompanied by a related question and answer pair in natural language. We utilize two datasets from chart question answering: ChartQA Masry et al. (2022) and PlotQA Methani et al. (2020a). To adapt these datasets for ChartVE, we apply a QA2Claim model Huang et al. (2023b) to transform the question-answer pairs into declarative statements. Paired with the original charts, these statements form positive instances (Entailment).
Meanwhile, chart captioning datasets include charts accompanied by descriptive captions, which typically consist of multiple sentences. Repurposing captions from VisText Tang et al. (2023a) and Chart-To-Text Kantharaj et al. (2022), we segment them into individual sentences. Each sentence is then paired with the relevant chart to create a positive instance. These adaptations allow us to repurpose existing datasets to create positive training examples for ChartVE.
Table-guided Negative Data Generation
Creating negative training samples is a vital step for balancing our dataset, which we achieve by perturbing the positive instances grounded in the underlying data tables of the charts. For any given chart $\mathcal{E}_{i}$, we utilize its source data table $\mathcal{A}_{{\mathcal{E}_{i}}}$ as the basis for generating factual inconsistencies. Our approach involves scanning through each value within $\mathcal{A}_{{\mathcal{E}_{i}}}$ and identifying if it appears as a substring within the corresponding positive caption sentence $c^{+}_{i}$. When a match is found, we introduce a factual error by substituting that substring with a different value from the same column in $\mathcal{A}_{{\mathcal{E}_{i}}}$, yielding a value or label-error infused negative sentence $c^{-}_{i}$. This approach ensures $c^{-}_{i}$ is relevant but factually inconsistent to the corresponding chart $\mathcal{E}_{i}$.
To incorporate trend errors, we rely on a predefined lexicon of trend-specific terminology, such as words denoting “increase” or “decrease.” We then detect these terms within $c^{+}_{i}$ and systematically replace them with their antonyms to generate a trend-dissenting statement.
Additionally, we synthesize out-of-context errors by pairing chart $\mathcal{E}_{i}$ with an unrelated caption sentence $c^{+}_{j}$ from the pool, ensuring that $c^{+}_{j}$ corresponds to a different chart, where $i\neq j$. This form of error generation effectively emulates captions that introduce concepts or data not represented in the target chart, broadening the set of negative samples for our training corpus.
The above process is illustrated in Algorithm 1. We use the training, development, and test sets of the repurposed datasets for training, validating, and testing ChartVE. This is vital for ensuring that ChartVE is free from data contamination in downstream applications. In total, we collected over 595K instances partitioned into training, development, and test splits with a ratio of 522:36:37, respectively.
Learning ChartVE
We selected UniChart as our base model, given its superior performance amongst comparable-size models [^5] Our fine-tuning begins with this checkpoint: https://huggingface.co/ahmed-masry/unichart-base-960.. Recognizing that UniChart has been pre-trained on chart question answering tasks, we employ a tailored input template $t$ as follows:
Does the image entail this statement: “SENTENCE”?
In this template, SENTENCE replaces the chart caption sentence $c$. Taking in a chart $\mathcal{E}$ and template $t$ as input, UniChart is fine-tuned to produce the token “yes” if the chart $\mathcal{E}$ entails the caption sentence $c$, and “no” otherwise using maximum likelihood estimate. During inference time, we use the same input format and probe the logits corresponding to the “yes” ($l_{\text{yes}}$) and “no” ($l_{\text{no}}$) decoder tokens. Following this, we apply the softmax function to convert these logits into an entailment score $s(\mathcal{E},c)$ that ranges from 0 to 1:
$s(\mathcal{E},c)=\frac{e^{l_{\text{yes}}}}{e^{l_{\text{yes}}}+e^{l_{\text{no}}}}.$ (1)
Here, $e$ is the base of the natural logarithm. Finally, we compute the minimum of the entailment scores for all sentences within a caption, denoted by $S(\mathcal{E},\mathcal{C})$, where $\mathcal{C}$ represents the set of all caption sentences for chart $\mathcal{E}$:
$S(\mathcal{E},\mathcal{C})=\min_{c\in\mathcal{C}}s(\mathcal{E},c).$ (2)
Meta-evaluation of Different Evaluation Metrics
Chocolate Model Lvlm Llm Ft LLaVA-1.5-13B 0.002 0.057 0.214 Bard -0.014 0.105 0.291 GPT-4V 0.157 0.205 0.215 DePlot + GPT-4 0.129 0.117 0.109 ChartVE (Ours) 0.178 0.091 0.215
Table 3: Kendall’s Tau correlation of different approaches on the Chocolate dataset.
To evaluate the effectiveness of different methods in assessing the factuality of generated captions on the Chocolate dataset, we employ Kendall’s Tau Kendall (1938) to compute the correlation between these methods and human judgments. Given the absence of prior work specifically focused on building factual inconsistency detection models for chart captions, we compare our ChartVE metric with methods that are capable of performing zero-shot reasoning including DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a), Bard Google (2023a), GPT-4V OpenAI (2023b), and the leading open-source LVLM, LLaVA-1.5-13B Liu et al. (2023c). The prompts for using these models as evaluation metrics can be found in Appendix E.
The meta-evaluation results are summarized in Table 3. We observed that, overall, metrics exhibit the strongest correlation with human judgment on the Ft split and the weakest on the Lvlm split. This pattern aligns with expectations: the Ft captions are littered with more obvious mistakes, including a substantial number of out-of-context and nonsense errors, as detailed in Figure 1. Conversely, errors stemming from LVLMs often demand intricate inferences regarding the data points’ positions relative to the axes, thus presenting a more substantial challenge for detection. Importantly, our ChartVE metric demonstrates the highest correlation on the most demanding Lvlm split. However, ChartVE underperforms on the Llm split, potentially due to shifts in token distribution, as DePlot + GPT-4 occasionally employs table-centric terminology (e.g., “columns” and “entries”) absent from ChartVE’s training data. Despite this, ChartVE compares favorably to proprietary LVLMs and, notably, outstrips the performance of the SOTA open-source LVLM, LLaVA-1.5-13B, despite ChartVE being 100 times smaller in scale.
On the Llm and Ft splits, the leading metrics are Bard and GPT-4V, respectively. Yet, what stands out is that Bard presents a negative correlation on the Lvlm split, signaling that LVLMs may falter significantly in evaluating the factuality of chart captions under certain conditions. As a result, we advocate for leveraging the best-performing metric for each split for evaluation.
4 Methodology
Figure 2: An overview of C2TFec. Our approach decomposes visual reasoning into image-to-structure rendering and text-based reasoning, allowing for interpretability and better correction of chart captions.
To address the factual errors identified in generated captions, we propose a framework, C2TFec, that capitalizes a two-step correction process that offers interpretability. C2TFec first transforms input charts into structured data tables (Section 4.1) and then refines captions by rectifying any discrepancies using the tabular data (4.2). Our deep dive into the performance of a joint framework, “DePlot + GPT-4”, revealed that a notable proportion of errors in caption generation originated from the DePlot component. To mitigate this, we have developed a superior chart-to-table conversion model that builds upon a stronger base model, UniChart, vastly fortified with expansive fine-tuning datasets. The primary advantage of this refined methodology is its ability to harness the strong reasoning strengths of GPT-4 to faithfully correct errors, thereby significantly upscaling the factuality of the generated captions. [^6] Here, we do not consider approaches based on LVLMs due to their tendency towards factual errors as evidenced in Figure 1.
[…]
Table 4: Correction performance of different models on the Chocolate dataset. ChartVE measures factuality by computing the entailment probability from each chart to the corresponding caption sentences. GPT-4V and Bard, when used as evaluation metrics, rate each chart caption as factually consistent with the chart or not. Levenshtein computes the edit distance between the corrected caption and the original caption (denoted as “N/A”). Metrics scores are shown separately for each of the three data splits based on captioning model source. The highest and second highest performing models per evaluation metric and split are highlighted in boldface and underlines respectively.
4.1 Chart-To-Table Conversion
The training data for our chart-to-table model is sourced from datasets including VisText Tang et al. (2023a), Chart-to-text Kantharaj et al. (2022), ChartQA Masry et al. (2022), and PlotQA Methani et al. (2020b), where we repurpose original charts and underlying data tables for our model’s training. We collected a total of 65K instances with a train:dev:test split of 61:2:2. In alignment with DePlot Liu et al. (2023a), our model is also trained to generate chart titles, enhancing its ability to contextualize the data represented in table form. Let $\mathcal{M}$ denote our proposed model. For a given chart figure $\mathcal{E}$, the model autoregressively generates a chart title $\mathcal{T}$ and a corresponding table $\mathcal{A}$ (i.e. $\mathcal{T},\mathcal{A}=\mathcal{M}(\mathcal{E})$).
4.2 Table-based Error Rectification
With the input chart now converted into structured tabular data, the second phase targets identifying and fixing factual inconsistencies in the initial caption $\mathcal{C}$. Our approach exploits the reasoning capacity of large language models, using GPT-4 to logically analyze the alignment between $\mathcal{C}$ and the constructed table $\mathcal{A}$. Specifically, GPT-4 first provides an explanatory breakdown of detected factual errors in $\mathcal{C}$ based on the table contents. It then uses this explanation to produce an edited caption $\mathcal{\hat{C}}$ remedying the identified inconsistencies. This transparent process enables users to validate the reasoning behind each correction.
Overall, the proposed C2TFec framework aims to produce verified chart captions by systematically isolating the language generation component from the analytical factual verification, taking advantage of the complementary strengths of separate vision and language models tailored to their respective domains. We hypothesize that the intermediate symbolic representation can act as an effective bridge to enforce and validate factual consistency.
5 Experimental Settings
To evaluate our proposed C2TFec for factual error correction in chart captions, we conduct experiments on our annotated Chocolate dataset.
Datasets
We utilize the 1,187 chart-caption pairs from the Chocolate dataset, comprising charts paired with both factually consistent and inconsistent captions generated by various models. The captions contain a diverse range of factual errors annotated by human judges, allowing comprehensive assessment. As described in Section 2, the dataset is divided into Lvlm, Llm, and Ft splits based on the caption models.
Baselines
Since Chocolate does not comprise of training data, we compare C2TFec against competitive baselines that are capable of performing strong zero-shot reasoning. This includes state-of-the-art open-source and proprietary LVLMs, LLaVA-1.5-13B Liu et al. (2023c), GPT-4V OpenAI (2023b), and Bard Google (2023a), as well as a LLM-based pipeline, DePlot + GPT-4 Liu et al. (2023a); OpenAI (2023a). For a fairer comparison between our approach and DePlot, we continue fine-tuning DePlot for an additional 5,000 steps on VisText, an approach which has been shown effective for adapting models to unseen domains Huang et al. (2023b). We denote this model as $\text{DePlot}_{\text{CFT}}$. The prompts used for each model are described in Appendix E.
Evaluation Metrics
We assess the factual consistency between corrected captions and input charts using the ChartVE metric, as well as GPT-4V and Bard, according to our recommendations in Section 3.2. In addition, since we require corrections to be made with as few edits as possible, we measure the number of edits using the Levenshtein distance Levenshtein et al. (1966).
6 Results
6.1 Main Results
The main results presented in Table 4 demonstrate that our proposed C2TFec framework achieves the best performance in terms of factual consistency on two out of three splits of the Chocolate dataset. Specifically, Chocolate attains the highest scores on the Lvlm and Ft splits, and achieves the second-best performance on the LLM. This indicates that the two-step process of first transforming charts into structured data tables and then rectifying factual inconsistencies using table-caption alignment is an effective strategy.
An interesting observation is that Bard struggles on the Lvlm split. Combined with the negative correlation with human judgments of factuality that we observed in Table 3, we can conclude that Bard is unreliable in identifying factual errors in chart captions. Additionally, when used as an evaluator, GPT-4V tends to assign unduly positive factuality scores to its own corrected outputs. On the Lvlm and Ft splits, GPT-4V rates its own performance as significantly higher than all other models (see Table 6), while other metrics show GPT-4V lagging behind C2TFec. This suggests GPT-4V may suffer from the self-enhancement bias Zheng et al. (2023), failing to reliably identify factual errors in its own generations when served as an evaluation metric. Hence, to further verify the effectiveness of our proposed approach, we conduct human evaluations in Section 6.2.
6.2 Human Evaluation
Our human evaluations aim to compare the performance of C2TFec and GPT-4V. The annotation tasks for human evaluation are the same as those presented to the annotators in Section 2: identifying various types of factual errors. Hence, the evaluation are conducted by the same annotators that perform the error analysis in Section 2. We sampled 30 charts from each split of Lvlm, Llm, and Ft. For each chart, human judges are presented with a caption generated by one of the models.
Figure 3 encapsulates the conclusive edge of C2TFec over GPT-4V across every error category. Notably, C2TFec demonstrates a marked reduction in Value Errors, surpassing GPT-4V by over a 20% margin within the Lvlm and Llm splits, and also half the overall error rate. Equally impressive was the near elimination of Trend Errors post-correction with C2TFec This underscores C2TFecs robust capability for accurate error rectification pertaining to reasoning regarding axes, including Label, Value, and Trend errors. A qualitative example is shown in Figure 4 to provide intuition on the superior correction ability of C2TFec over GPT-4V. We have also observed that the reason GPT-4V cannot correct these types of errors well is due to its inability to deduce the values of the data points from the input chart (see Figure 7).
In contrast, GPT-4V exhibited a comparative proficiency in addressing Out-of-context Errors, involving information extraneous to the chart’s scope. However, GPT-4V seemed challenged in rectifying errors within captions generated by itself, particularly within the Lvlm split. This observation echoes recent findings on LLMs’ inability to self-correct Huang et al. (2023a); Valmeekam et al. (2023), we find that LVLMs also cannot perform self-correction. More importantly, our human evaluation results, combined with our findings in Table 4 and Table 6, reflect that GPT-4V is subject to serious self-enhancement bias. Consequently, although GPT-4V’s capabilities are formidable, we recommend not using them to assess their own outputs.
Figure 3: Human evaluation results on subsets of the Chocolate dataset, comparing C2TFec and GPT-4V. C2TFec corrects significantly more errors compared to GPT-4V, especially Value, Label, and Trend Errors.
Figure 4: An example showing how decomposing the visual reasoning process into image-to-structure rendering and text-based reasoning allows C2TFec to accurately rectify errors in chart captions. Texts marked in red indicate non-factual information units in the caption, whereas those marked in blue represent information units faithful to the chart. In this instance, C2TFec successfully corrects all Value and Label Errors presented in the original caption. Conversely, GPT-4V fails to identify the factual inconsistencies and merely reorders the entities in the caption.
7 Related Work
7.1 Chart Captioning
Chart captioning is essential for accurately interpreting and communicating the information conveyed by chart images, particularly in news articles and social media, where factuality is imperative to prevent misinformation. While current datasets like FigureQA Kahou et al. (2017), DVQA Kafle et al. (2018), PlotQA Methani et al. (2020c), VisText Tang et al. (2023a), and Chart-to-Text Kantharaj et al. (2022) offer chart image descriptions and question-answer pairs to train models, advancements in vision-language models like ChartT5 Zhou et al. (2023), MatCha Liu et al. (2023b), and UniChart Masry et al. (2023) have largely prioritized relevance and fluency over factual accuracy. Our work provides a rigorous characterization of factual errors in chart captioning and comparisons of methods to address this gap. By focusing on faithfulness and correction, we complement the emphasis of prior work and aim to produce more trustworthy chart captions.
7.2 Factual Error Correction
Previous research on based factual error correction has focused on two textual tasks: text summarization and fact-checking. Within text summarization, the bulk of work, such as Dong et al. (2020), has been carried out in the news domain and often involves methods that substitute inconsistent entities from the source text. Some studies have enhanced this approach through entity-replacement reranking techniques Chen et al. (2021), autoregressive models for rewriting and perturbation filtering Cao et al. (2020); Zhu et al. (2021); Adams et al. (2022), and editing strategies that focus on selective deletion Wan and Bansal (2022). In contrast, Fabbri et al. (2022) employed sentence compression datasets to train their models. More recently, Gao et al. (2023) have expanded the focus of these studies to include dialogue summarization.
Moving to the domain of fact-checking, this area has experienced a flurry of activity, particularly with the increased attention on combating misinformation Fung et al. (2021); Wu et al. (2022); Fung et al. (2022); Huang et al. (2023d, c); Qiu et al. (2023). Early approaches train a distantly supervised model that involves a masker and a corrector Shah et al. (2020); Thorne and Vlachos (2021). Thorne and Vlachos (2021) made significant strides by developing the first factual error correction dataset for fact-checking, thus enabling fully supervised training for error correctors. Recently, Huang et al. (2023b) propose an interpretable framework that breaks down the process of fact-checking into individual components. Our study builds on these insights and extends them to a multimodal context, which challenges models to understand the chart images and the consistency between different modalities.
8 Conclusion
Our study exposes the prevalent issue of factual errors in chart captions generated by LVLMs, such as GPT-4V, and introduces Chocolate to scrutinize these errors. We establish the Chart Caption Factual Error Correction task to propel the creation of trustworthy captioning systems and present ChartVE, an evaluation model surpassing LVLMs in mirroring human assessments of caption factuality. Our two-stage correction framework, C2TFec, provides an interpretable means of improving caption factuality by transforming visual data into structured tables for more faithful error corrections. Our work marks an essential step in ensuring verifiable and trustworthy chart captions. Future directions include extending our approach to multimodal contexts beyond charts, developing more sophisticated error detection and correction algorithms, and creating datasets covering a broader range of visual content.
9 Limitations
We acknowledge that our study did not rigorously examine the sensitivity of different systems to the variations in the prompts used. The effectiveness of several natural language processing tasks is known to be influenced by the design of the input prompts. Our omission of a systematic sensitivity analysis means that there could be a range of responses to different prompts that we have not accounted for, which may affect the generalization of our results. However, we did not perform prompt tuning to craft prompts that benefit our proposed model. Therefore, the comparisons across all models are fair. Due to the scope of our study, we leave the prompt sensitivity experiments for future work.
In addition, most charts in the datasets we used are line plots and bar plots. Future efforts can extend our work with additional analyses for other types of charts, such as scatter plots, violin plots, and distribution plots.
Appendix A Further Discussions
Captioning Model Analysis
In addition the findings we summarize in Section 2.2, we also found that the error distribution for each model differs on different datasets. Almost all models make significantly more Nonsense Errors on the Pew dataset. In addition, task-specific models observe a non-negligible increase in Out-of-context Errors on the Pew dataset. Both observations could be explained by the fact that these models are sometimes confused about the charts in Pew, which are often associated with more complicated structures.
Main Results
We see that C2TFec outperforms the pipeline approaches of DePlot/$\text{DePlot}_{\text{CFT}}$ + GPT-4 across the board. While both methods utilize an intermediate tabular representation and leverage GPT-4 for language generation/correction, C2TFec employs a superior chart-to-table conversion model with much more comprehensive training datasets. This results in extracted tables that more faithfully capture the underlying chart data, better facilitating the downstream factual error correction. C2TFec also requires a relatively small number of edits to captions according to Levenshtein distance, making focused changes to improve factuality while minimizing revisions. An example output from C2TFec is shown in Figure 4. By comparison, the proprietary LVLM Bard produces corrected captions requiring 127.83 as many character-level edits on average. This signals excessive rewriting rather than targeted error correction.
Appendix B Annotation Details
In this section, we present the details of our human annotation conducted in Section 2.
B.1 Worker Qualification
We laid out specific preliminary criteria for the recruitment of MTurk workers with impressive performance records. These prerequisites comprise a HIT approval percentage of 99% or above, a minimum of 10,000 approved HITs, and the worker’s location within the United Kingdom, Canada, or the United States.
Moreover, beyond these initial criteria, suitable workers have to successfully pass two staged qualification examinations focused on identifying factual errors in generated chart captions. To optimize the qualification procedure, the authors manually annotate two HITs, each consisting of one chart and one caption produced by one of our chart captioning models. In every qualification round, annotators are exposed to one of these annotated examples. Workers whose annotations fail to correspond closely with ours are eliminated from the selection procedure.
Finally, a group of 7 annotators who successfully navigated all three stages of qualification tests were chosen. Additionally, each HIT was meticulously crafted to ensure that annotators could achieve an equivalent hourly pay rate of \$15 - \$20, assuming they work without interruption.
B.2 Annotation Guidelines
In this task, you will evaluate the factual errors for a generated caption with regard to the reference chart. To correctly solve this task, follow these steps:
  • Carefully read the generated caption and the reference chart.
  • Compare the generated caption against the reference chart and decide whether the caption contains any factual error defined below.
  • You should click/press the button if an error occurs. A blue button indicates the caption contains the corresponding factual error, while a white button means the caption does not contain such an error.
Warning: Annotations will be checked for quality against control labels, low-quality work will be rejected.
Error definition
  • Value error: A quantitative data value is incorrect.
  • Label error: A non-quantitative data value is incorrect.
  • Trend error: The direction of a trend is wrong.
  • Magnitude error: The magnitude or variance of a trend is wrong.
  • Out-of-context error: The caption introduces concepts that are not present in the chart.
  • Grammatical error: The grammar of the caption is wrong.
  • Nonsense error: The caption is incomplete or does not make sense at all.
B.3 Annotation Interface
The interface for our human annotation is shown in Figure 5.
Appendix C Dataset Details
Table 5 presents the detailed statistics of each split in our dataset.
Chocolate-Lvlm Chocolate-Llm Chocolate-Ft # Factual # Non-factual # Factual # Non-factual # Factual # Non-factual Sentence 1,683 1,270 518 469 360 1,023 Caption 74 321 27 169 112 484
Table 5: Dataset statistics per pplit. A sentence is considered factual if and only if it does not contain any factual error. A caption is considered factual if all its sentences are factual.
Appendix D Implementation Details
Input: Data table $\mathcal{A}_{\mathcal{E}_{i}}$ for chart $\mathcal{E}_{i}$, Positive caption sentence $c^{+}_{i}$.
Output: Set of negative caption sentences $C^{-}_{i}=\{c^{-}_{i,\text{value}},c^{-}_{i,\text{trend}},c^{-}_{i,\text{context}}\}$.
1 Initialize $C^{-}_{i}$ as an empty set;
2 Define a lexicon of trend terms $T$;
3 Define entailment threshold $\tau$;
4 // Generate Value and Label Errors;
5 for each cell value $v$ in $\mathcal{A}_{\mathcal{E}_{i}}$ do
6       if $v$ is a substring of $c^{+}_{i}$ then
7             Randomly sample a new value $v^{\prime}$ from the same column in $\mathcal{A}_{\mathcal{E}_{i}}$;
8             Replace $v$ in $c^{+}_{i}$ with $v^{\prime}$ to get $c^{-}_{i,\text{value}}$;
9             Add $c^{-}_{i,\text{value}}$ to $C^{-}_{i}$;
10
11
12// Generate Trend Errors;
13 for each trend term $t$ in $T$ do
14       if $t$ is found in $c^{+}_{i}$ then
15             Replace $t$ in $c^{+}_{i}$ with its antonym to get $c^{-}_{i,\text{trend}}$;
16             Add $c^{-}_{i,\text{trend}}$ to $C^{-}_{i}$;
17
18
19// Generate Out-of-Context Errors;
20 Randomly select a different chart $\mathcal{E}_{j}$ where $j\neq i$;
21 Pair $\mathcal{E}_{i}$ with unrelated caption sentence $c^{+}_{j}$ to get $c^{-}_{i,\text{context}}$;
22 Add $c^{-}_{i,\text{context}}$ to $C^{-}_{i}$;
23 return $C^{-}_{i}$;
Algorithm 1 Table-guided Negative Data Generation
D.1 Details of the Chart-To-Table Model
[…]
D.2 Table-guided Negative Data Generation
In Algorithm 1, we depict the details of how we generate negative data for our ChartVE model.
D.3 Model Training
The Chart-To-Table model and ChartVE are optimized using AdamW for a maximum of 20,000 and 50,000 steps, respectively. The learning rates for both models are set to 5e-5. During inference time, the Chart-To-Table model uses beam search with a beam width of 4.
[…]
Table 6: Correction performance on Chocolate using GPT-4V as the evaluation metric. GPT-4V, when used as an evaluator, assigns significantly higher scores to its own generations. This suggests potential self-enhancement bias of GPT-4V. Note that GPT-4V also assign a high scores to the original captions (i.e. N/A) on the Lvlm split. This is because half of these captions are directly generated from GPT-4V.
Figure 6: The impact of value labeling. We prompted GPT-4V to generate captions of two charts we created using the Seaborn library from an underlying table sampled from the Chart-to-Text dataset, with or without labeling the values of the bars on the chart. We see that when the labeled values are presented in the chart, GPT-4V is capable of producing more factual captions.
Figure 7: An example showing GPT-4V cannot accurately extract tables from charts. This indicates its inability to infer the actual value of each data point within the chart.
Appendix E Prompts Used In Experiments
The prompts for using LVLM and LLM as evaluation metrics are displayed in Figure 8 and Figure 9, while the prompts for factual error correction are shown in Figure 10 and Figure 11.
Figure 8: Prompts for using GPT-4V, Bard, and LLaVA-1.5 as a evaluator.
Figure 9: Prompts for using DePlot + GPT-4 as a evaluator.
Figure 10: Prompts for using GPT-4V, Bard, and LLaVA as a factual error corrector.
Figure 11: Prompts used for using DePlot + GPT-4 as a factual error corrector.

### v2

@@
 Abstract
-Recent advancements in large vision-language models (LVLMs) have led to significant progress in generating natural language descriptions for visual content and thus enhancing various applications.
-One issue with these powerful models is that they sometimes produce texts that are factually inconsistent with the visual input.
-While there has been some effort to mitigate such inconsistencies in natural image captioning, the factuality of generated captions for structured document images, such as charts, has not received as much scrutiny, posing a potential threat to information reliability in critical applications.
-This work delves into the factuality aspect by introducing a comprehensive typology of factual errors in generated chart captions.
-A large-scale human annotation effort provides insight into the error patterns and frequencies in captions crafted by various chart captioning models, ultimately forming the foundation of a novel dataset, Chocolate.
-Our analysis reveals that even state-of-the-art models, including GPT-4V, frequently produce captions laced with factual inaccuracies.
-In response to this challenge, we establish the new task of Chart Caption Factual Error Correction and introduce ChartVE, a model for visual entailment that outperforms proprietary and open-source LVLMs in evaluating factual consistency.
+Advances in large vision-language models (LVLMs) have led to significant progress in generating natural language descriptions for visual contents.
+These powerful models are known for producing texts that are factually inconsistent with the visual input.
+While some efforts mitigate such inconsistencies in natural image captioning, the factuality of generated captions for structured visuals, such as charts, has not received as much scrutiny.
+This work introduces a comprehensive typology of factual errors in generated chart captions.
+A large-scale human annotation effort provides insight into the error patterns in captions generated by various models, ultimately forming the foundation of a dataset, Chocolate.
+Our analysis reveals that even advanced models like GPT-4V frequently produce captions laced with factual inaccuracies.
+To combat this, we establish the task of Chart Caption Factual Error Correction and introduce ChartVE, a visual entailment model that outperforms current LVLMs in evaluating caption factuality.
 Furthermore, we propose C2TFec, an interpretable two-stage framework that excels at correcting factual errors.
-This work inaugurates a new domain in factual error correction for chart captions, presenting a novel evaluation mechanism, and demonstrating an effective approach to ensuring the factuality of generated chart captions.
-[^1] The code and data will be made publicly available: https://github.com/khuangaf/CHOCOLATE.
+This work inaugurates a new domain in factual error correction for chart captions, presenting a novel evaluation metric, and demonstrating an effective approach to ensuring the factuality of generated chart captions.
+The code and data as well as the continuously updated benchmark can be found at: https://khuangaf.github.io/CHOCOLATE/.
 Do LVLMs Understand Charts?
 Analyzing and Correcting Factual Errors in Chart Captioning
-[…]
+Kung-Hsiang Huang1,2∗    Mingyang Zhou3     Hou Pong Chan4∗     Yi R.
+Fung1     Zhenhailong Wang1     Lingyu Zhang3     Shih-Fu Chang3     Heng Ji1 1University of Illinois Urbana-Champaign     2Salesforce AI Research 3Columbia University     4DAMO Academy, Alibaba Group 1{yifung2, wangz3, hengji}@illinois.edu   2kh.huang@salesforce.com 3{mz2974, lz2814, sc250}@columbia.edu   4houpong.chan@alibaba-inc.com
+**footnotetext: Work was done while Kung-Hsiang was at UIUC and Hou Pong was at the University of Macau.
 1 Introduction
-Large vision-language models (LVLMs) have recently shown impressive capabilities in generating natural language descriptions of visual content like images, videos and charts Li et al.
-(2020, 2022); Wang et al.
-(2022); OpenAI (2023b); Google (2023a); Liu et al.
+Large vision-language models (LVLMs) have recently shown impressive capabilities in generating natural language descriptions of visual content like images, videos and charts OpenAI (2023b); Google (2023a); Liu et al.
 (2023c); Wang et al.
@@
 Chart captioning is particularly important for data analysts, business analysts, and journalists who rely on accurate chart interpretations for decision-making and reporting.
-However, no prior work have studied the factuality of the generated captions.
-Factuality, also known as the faithfulness or factual consistency between inputs and outputs, of generated captions is of critical importance when leveraging LVLMs for both practical applications and research advancements.
-In the domain of information dissemination, for instance, such models can be deployed to automatically generate descriptions for charts in news articles Liu et al.
-(2021), educational materials Fu et al.
-(2022), and on social media platforms Monteiro et al.
-(2017).
-Ensuring that these captions are factual directly impacts the credibility and reliability of the content provider.
-To better understand the factual errors produced by chart captioning models, we first introduce a typology of factual errors particularly relevant in the chart domain.
-Using this scheme, we conduct a large-scale human annotation study to analyze the distributions and frequencies of various error types, such as Value Error and Label Error, across captions produced by various chart captioning models, from task-specific fine-tuned models to LVLMs (see Table 1).
-The annotated samples are then categorized into three splits, Lvlm (Large-vision Language Models), Llm (Large Language Models), and Ft (Fine-tuned), based on the architecture and the scale of the underlying models that generate the captions, and form a dataset which we named Chocolate.
+However, no prior work has studied the factuality [^1] Factuality is also known as the faithfulness or factual consistency between inputs and outputs of the generated captions.
+Given that factuality is vital for credibility in applications of chart captioning in news articles Liu et al.
+(2021), educational resources Fu et al.
+(2022), and social media Monteiro et al.
+(2017), examining the truthfulness of generated captions is a critical concern.
+To understand the factual errors in chart captioning models, we introduce a typology of factual errors for the chart domain.
+Using this scheme, we conduct a large-scale human annotation study to analyze the distributions of various error types, such as Value Error and Label Error, in captions from various models, from task-specific fine-tuned models to LVLMs (see Table 1).
+The annotated samples are then categorized into three splits, Lvlm (Large-vision Language Models), Llm (Large Language Models), and Ft (Fine-tuned Vision-language Models), based on the architecture and the scale of the underlying models, and form a dataset which we named Chocolate.
 With this dataset collected, we aim to answer three main research questions.
@@
 We find the answer is no (Section 2).
-Out of the 1,187 captions we analyzed, only 213 of them are factual (see Table 2).
-Even state-of-the-art LVLM, GPT-4V OpenAI (2023b), produces a great portion of errors in its generated captions (see Figure 1).
-In addition, we show that different chart captioning models exhibit substantially different factual error distributions in generated captions across different chart domains.
-The prevalence of factual inconsistencies observed in the generated captions by various models underscores the urgent need to address the factual accuracy of such technologies.
-Consequently, we introduce a task named Chart Caption Factual Error Correction (Section 3), which presents a novel challenge of rectifying factual inaccuracies in chart captions generated by LVLMs.
-A pertinent question that arises from this task is: how can we automatically evaluate the factual consistency between charts and captions/corrections?
-To tackle this question, we conceptualize the evaluation procedure as a task of visual entailment.
-We present the first model for chart visual entailment, named ChartVE.
-This model is trained by repurposing existing datasets from chart summarization and chart question answering.
-Our comprehensive correlation analysis shows that ChartVE performs competitively with proprietary LVLMs, GPT-4 and Bard Google (2023a), and outperforms an open-source LVLM, LLaVA-1.5 Liu et al.
-(2023c), despite being considerably smaller—100 times less in size.
-The findings emphasize the efficiency and efficacy of ChartVEn evaluating factual accuracy for the task of Chart Caption Factual Error Correction.
+Specifically, 82.06% of the generated captions are non-factual (see Table 2).
+Even state-of-the-art LVLMs like GPT-4V OpenAI (2023b) produce a great portion of errors in its generated captions (see Figure 1).
+The prevalence of factual inconsistencies observed in the generated captions by various models underscores the urgent need to mitigate the factual errors of such models.
+Hence, we introduce a new task, Chart Caption Factual Error Correction (Section 3), which presents a novel challenge of rectifying factual inaccuracies in chart captions generated by LVLMs.
+A pertinent question that arises from this task is: how to automatically evaluate the factual consistency between charts and captions?
+To tackle this question, we present ChartVE, novel visual entailment approach to assess the factual consistency of chart captions.
+This model is trained by repurposing existing resources from chart summarization and chart question answering.
+Results show that ChartVE performs competitively with proprietary LVLMs and outperforms the most advanced open-source LVLM, despite being 64 times less in size.
 Figure 1: Error distribution for different models on VisText and Pew.
@@
 An error rate of 0.4 indicates that 40% of the sentences in the generated captions contain such an error.
-Note that a single caption may contain multiple error types; hence, the maximum value for a stacked bar is greater than 1.0.
-We show that even the most advanced LVLM, GPT-4, generates captions with a high rate of factual error.
+Note that a single caption may contain multiple types of errors; hence, the maximum value for a stacked bar is greater than 1.0.
+We show that even the most advanced LVLM, GPT-4V, generates captions with a high rate of factual error.
 Now that we have set up the task, we turn to the challenge of how to effectively correct factual errors in chart captions?
-Here, we propose C2TFec (Section 4), an interpretable two-step framework that breaks down visual reasoning into image-to-structure rendering and text-based reasoning.
+We propose C2TFec (Section 4), an interpretable two-step framework that decomposes visual reasoning into image-to-structure rendering and text-based reasoning.
 C2TFec first transforms the input chart into a structured data table representation.
-Grounded in this extracted tabular data, the second component then identifies and fixes any factual inconsistencies in the initial caption through a transparent reasoning process.
+Grounded in this extracted tabular data, the second component then identifies and fixes any factual inconsistencies in the generated caption through an interpretable reasoning process.
 Our experiments demonstrate that this explicit decomposition enables more reliable factuality corrections compared to end-to-end approaches.
-The intermediate symbolic data acts as an effective bridge between charts and captions.
-By exploiting the complementary strengths of a specialized Chart-To-Table model and GPT-4, C2TFec significantly outperforms competitive baselines including GPT-4V, as validated through both automatic evaluation and human assessment.
+The intermediate symbolic representation acts as an effective bridge between charts and captions, enabling C2TFec to significantly outperform competitive baselines including GPT-4V (Section 6).
 In summary, our contributions are as follows:
-  • We present the first empirical analysis of factual errors in captions produced by models of various scales and architectures.
-  • We introduce the Chart Caption Factual Error Correction task that challenges models to correct factual errors in generated chart captions and construct the Chocolate dataset for this task by defining a novel typology of factual errors in chart captions.
-  • We present ChartVE, a reference-free evaluation metric based on visual entailment that correlates better with human judges compared to LVLM-based approaches.
-  • We propose C2TFec, an interpretable two-stage error correction framework that demonstrates better performance compared to all existing LVLMs.
-2 Dataset Creation
+  • We present the first analysis of factual errors in captions produced by models of various scales using a novel error typology, which results in the Chocolate dataset.
+  • We introduce the Chart Caption Factual Error Correction task that challenges models to correct factual errors in generated chart captions.
+  • We present ChartVE, a reference-free evaluation metric based on visual entailment that correlates better with human judges than LVLMs.
+  • We propose C2TFec, an interpretable two-stage error correction framework that performs better than all existing LVLMs.
+2 Analyzing Factual Errors
 To understand the capabilities of existing models in summarizing key information from charts, we conduct a large-scale analysis on six most advanced chart captioning models on the VisText Tang et al.
@@
 To facilitate this process, we introduce an error typology, as illustrated in Section 2.1.
-Upon gathering human annotations, we discuss the quality of the collected data (Section 2.3) and present a detailed analysis of different captioning models (Section 2.2).
+Upon gathering human annotations, we present a detailed analysis of different captioning models (Section 2.2) and discuss the quality of the collected data (Section 2.3).
 2.1 Error Typology
@@
 Chart Category Example Caption
-Value Error
-Asians have a turnout rate of 20.4% in 1990.
-Label Error
-Asians have the highest turnout rates across the years.
-Trend Error
-From 1986-2014, the turnout rates are increasing overall.
-Magnitude Error
-From 1986-2014, the turnout rates are sharply decreasing overall.
-Out-of-context Error
-Vietnamese have the highest turnout rates among Asians.
-Nonsense Error
-From 1986-2014, #?sep #sep #sep #sep.
-Grammatical Error
-The turnout rates are decrease overall.
+[Uncaptioned image] Value Error Asians have a turnout rate of 20.4% in 1990.
+Label Error Asians have the highest turnout rates across the years.
+Trend Error From 1986-2014, the turnout rates are increasing overall.
+Magnitude Error From 1986-2014, the turnout rates are sharply decreasing overall.
+Out-of-context Error Vietnamese have the highest turnout rates among Asians.
+Nonsense Error From 1986-2014, #?sep #sep #sep #sep.
+Grammatical Error The turnout rates are decrease overall.
 Table 1: Typology of errors illustrated with an example chart.
@@
 Trend Error
-The overall direction of change over time or comparison between groups is incorrectly described in the caption, such as stating an increasing trend when the values in the chart are decreasing.
+The overall direction of change over time or comparison between groups is incorrectly described in the caption, such as stating an increasing trend when it is actually decreasing.
 Magnitude Error
-The degree or amount of difference described for a trend or comparison is factually inconsistent with the chart, such as stating a small/sharp increase/decrease when the change is more gradual based on the chart’s data.
+The degree or amount of difference described for a trend is unfaithful to the chart, such as stating an increase “sharp” when the chart shows it is actually “smooth”.
 Out-of-context Error
-Concepts, variables, or any information introduced in the caption that do not exist at all in the content of the chart.
+Concepts, variables, or any information introduced in the caption that does not exist at all in the content of the chart.
 The caption contains factual statements not grounded in the actual chart contents.
@@
 Grammatical Error
-There are grammatical mistakes in the structure or syntax of the caption.
-[^2] Note that we do not consider grammatical errors as factual inconsistency.
-We analyze grammatical errors to understand how fluent the captioning models are.
+There are grammatical mistakes in the structure or syntax of the caption. [^2] Note that we do not consider grammatical errors as factual inconsistency. They are analyzed for assessing fluency.
 2.2 Captioning Model Analysis
-We consider the various types of models.
+We consider various types of models.
 First, ChartT5 Zhou et al.
@@
 (2023a); OpenAI (2023a) is a LLM-based pipeline approach.
-Finally, GPT-4V OpenAI (2023b) and Bard Google (2023a) [^3] We tested Bard before the release of Gemini Google (2023b).
+Finally, GPT-4V and Bard [^3] We tested Bard before Gemini’s release Google (2023b).
 are the strongest LVLMs.
@@
 Invalid output sequences, such as empty strings, are filtered out.
-We compute the percentage of sentences with factual errors for different models and datasets, with a breakdown of the proportion of different error types.
-Here, we focus on the error distribution on the sentence level instead of the caption level since different models generate captions of different lengths.
-A sentence-level evaluation can help mitigate this discrepancy and facilitate a fairer comparison.
+We compute the percentage of sentences with factual errors for different models and datasets, with a breakdown of different error types.
+Error rates are computed at the sentence level instead of the caption level since different models generate captions of different lengths.
+A sentence-level evaluation helps mitigate this discrepancy and facilitates a fairer comparison.
 From Figure 1, we made the following observations.
 First, SOTA chart captioning models often fail to produce factual captions.
-Additionally, as shown in Table 2, we calculated the percentage of non-factual captions, revealing that 82.05% of captions contain at least one factual error.
-More importantly, even advanced models like GPT-4V and Bard, which have demonstrated proficiency in a variety of vision-language tasks, produce factually incorrect captions 81.27% of the time, as recorded in Table 5.
-These findings highlight the inherent difficulties of chart understanding tasks and the limitations of SOTA vision-language models.
+Additionally, as shown in Table 2, we calculated the percentage of non-factual captions, revealing that 82.06% of captions contain at least one factual error.
+More importantly, even models like GPT-4V and Bard, which have demonstrated proficiency in a variety of vision-language tasks, produce factually incorrect captions 81.27% of the time, as recorded in Table 10.
+These findings highlight the inherent difficulties of chart captioning tasks and the limitations of SOTA vision-language models.
 Second, task-specific chart captioning models and LVLMs show opposite trends on the two datasets.
-Task-specific fine-tuned models, including ChartT5, MatCha, and UniChart, produce fewer errors on the VisText Tang et al.
-(2023a) dataset.
+Task-specific models, including ChartT5, MatCha, and UniChart, produce fewer errors on the VisText dataset.
 Conversely, LVLMs, including GPT-4V and Bard, generate significantly fewer errors on the Pew dataset.
-The key distinctions on these datasets are two: (1) the vast majority of the charts in Pew datasets have corresponding numbers labeled on or next to each data point and (2) charts in VisText are associated with fewer complex structures compared to those in the Pew dataset.
-We hypothesize that LVLMs may be better at capturing the numbers labeled next to the data points in charts, whereas task-specific models are more capable of reading the exact value of the data points by aligning them to axes.
+The key distinctions on these datasets are two: (1) the prevalent labeled values on charts from Pew and (2) the simpler structures in charts from VisText.
+We hypothesize that LVLMs may be better at utilizing the labeled numbers, while task-specific effectively interpret values via axis alignment.
 We show an example to validate this hypothesis in Figure 6.
 Third, LVLMs cannot consistently outperform task-specific fine-tuned models.
-Although LVLMs are trained on data with a much larger scale using many more parameters, task-specific models may perform better than LVLMs when suitable pre-training objectives and model architectures are adopted.
-For example, on the VisText dataset, we see that UniChart achieves a much better performance than Bard and is comparable to GPT-4V in terms of producing more factual captions.
-This can be attributed to UniChart’s various pre-training objectives tailored for chart understanding, enabling it to better identify the relationship between data points within the charts from VisText.
-The dataset resulting from the analysis is named Chocolate (Captions Have Often Chosen Obfuscating Lies About The Evidence), where each instance consists of a chart, a generated chart caption, and error types labeled by human annotators.
+Despite their extensive training data and parameters, LVLMs may be surpassed by task-specific models with appropriate pre-training objectives and architectures.
+For example, on the VisText dataset, UniChart outperforms Bard and is comparable to GPT-4V in terms of producing more factual captions owing to UniChart’s various pre-training objectives for chart comprehension, enabling better interpretation of the relationship between data points within charts.
+The dataset resulting from the analysis is named Chocolate (Captions Have Often ChOsen Lies About The Evidence), where each instance consists of a chart, a generated chart caption, and error types labeled by human annotators.
 Drawing insights from Tang et al.
-(2023b) that factual errors produced by different kinds of models may be easier or more difficult to identify, we split Chocolate into three categories.
-First, the Lvlm split corresponds to captions produced by GPT-4V and Bard.
-Second, the Llm split contains captions generated by DePlot + GPT-4.
-Finally, the Ft split is associated with captions from ChartT5, UniChart, and MatCha.
-The details of each split are shown in Appendix C.
+(2023b) that factual errors produced by different kinds of models may be easier or more difficult to identify, we categorize Chocolate into three splits: the Lvlm split, with captions from GPT-4V and Bard; the Llm split, featuring DePlot + GPT-4 outputs; and the Ft split, for ChartT5, UniChart, and MatCha captions.
+Split details are in Appendix C.
 2.3 Dataset Quality
@@
 There was no established metric for evaluating the factual consistency between a chart and the corresponding chart caption.
-In addition, since our dataset does not contain annotated reference corrected captions [^4] We did not collect reference corrected captions due to the challenges of curating high-quality references through crowd-sourcing., text-based metrics such as SARI Xu et al.
-(2016) cannot be adopted.
+In addition, since our dataset does not contain annotated reference captions [^4] Reference captions are not collected due to the challenges of curating high-quality references through crowd-sourcing., text-based metrics cannot be adopted.
 As a solution, we propose ChartVE, a reference-free evaluation metric based on chart visual entailment, as detailed in the following paragraphs.
 ChartVE Overview
-We formulate the task of chart caption factual inconsistency detection as a chart visual entailment task.
-Given a chart caption sentence $c$ and a chart $\mathcal{E}$, the task is to predict whether the relationship from $\mathcal{E}$ to $c$ as Entailment (i.e.
-factually consistent) or NotEntailment (factually inconsistent).
+We formulate the inconsistency detection problem as a chart visual entailment task.
+Given a chart caption sentence $c$ and a chart $\mathcal{E}$, the task is to predict whether the relationship from $\mathcal{E}$ to $c$ as Entailment (factually consistent) or NotEntailment (factually inconsistent).
 The main challenge of learning a visual entailment model for this task is the lack of data.
-To overcome this challenge, we repurpose data from relevant tasks, such as chart QA and chart captioning, as positive samples.
+To overcome this challenge, we repurpose data from relevant tasks, such as chart QA, as positive samples.
 Then, we propose a table-guided negative data generation to produce negative samples.
-In the following paragraphs, we detail how each step is achieved.
 Positive Data Creation
 We consider datasets from two tasks that are closely related to the chart visual entailment task: chart question answering and chart captioning.
-The chart question answering task requires models to respond to questions regarding a specific chart, with each instance comprising a chart accompanied by a related question and answer pair in natural language.
 We utilize two datasets from chart question answering: ChartQA Masry et al.
 (2022) and PlotQA Methani et al.
-(2020a).
-To adapt these datasets for ChartVE, we apply a QA2Claim model Huang et al.
-(2023b) to transform the question-answer pairs into declarative statements.
-Paired with the original charts, these statements form positive instances (Entailment).
-Meanwhile, chart captioning datasets include charts accompanied by descriptive captions, which typically consist of multiple sentences.
-Repurposing captions from VisText Tang et al.
+(2020).
+Using a QA2Claim model Huang et al.
+(2023b), we transform the question-answer pairs into declarative statements and pair them with the original charts to form positive instances (Entailment).
+For chart captioning, captions from VisText Tang et al.
 (2023a) and Chart-to-Text Kantharaj et al.
-(2022), we segment them into individual sentences.
-Each sentence is then paired with the relevant chart to create a positive instance.
-These adaptations allow us to repurpose existing datasets to create positive training examples for ChartVE.
+(2022) are segmented into individual sentences.
+Each sentence is paired with the relevant chart to create a positive instance.
+These methods allow us to repurpose existing resources for training ChartVE.
 Table-guided Negative Data Generation
-Creating negative training samples is a vital step for balancing our dataset, which we achieve by perturbing the positive instances grounded in the underlying data tables of the charts.
-For any given chart $\mathcal{E}_{i}$, we utilize its source data table $\mathcal{A}_{{\mathcal{E}_{i}}}$ as the basis for generating factual inconsistencies.
-Our approach involves scanning through each value within $\mathcal{A}_{{\mathcal{E}_{i}}}$ and identifying if it appears as a substring within the corresponding positive caption sentence $c^{+}_{i}$.
-When a match is found, we introduce a factual error by substituting that substring with a different value from the same column in $\mathcal{A}_{{\mathcal{E}_{i}}}$, yielding a value or label-error infused negative sentence $c^{-}_{i}$.
-This approach ensures $c^{-}_{i}$ is relevant but factually inconsistent to the corresponding chart $\mathcal{E}_{i}$.
-To incorporate trend errors, we rely on a predefined lexicon of trend-specific terminology, such as words denoting “increase” or “decrease.” We then detect these terms within $c^{+}_{i}$ and systematically replace them with their antonyms to generate a trend-dissenting statement.
-Additionally, we synthesize out-of-context errors by pairing chart $\mathcal{E}_{i}$ with an unrelated caption sentence $c^{+}_{j}$ from the pool, ensuring that $c^{+}_{j}$ corresponds to a different chart, where $i\neq j$.
-This form of error generation effectively emulates captions that introduce concepts or data not represented in the target chart, broadening the set of negative samples for our training corpus.
+Generating negative training samples is achieved by perturbing the positive instances grounded in the underlying data tables of the charts.
+For a chart $\mathcal{E}_{i}$ and its underlying data table $\mathcal{A}_{{\mathcal{E}_{i}}}$, we locate values in $\mathcal{A}_{{\mathcal{E}_{i}}}$ that matches a substring within the positive caption $c^{+}_{i}$.
+When a match is found, the substring in the caption is substituted with a different value from the same column in $\mathcal{A}_{{\mathcal{E}_{i}}}$, yielding a value or label-error infused negative sentence $c^{-}_{i}$, maintaining relevance while ensuring inconsistency with $\mathcal{E}_{i}$.
+For trend-related errors, we replace trend-terms found in $c^{+}_{i}$ with their opposites, drawing on a specific lexicon of terms like “increase” and “decrease,” thereby creating trend-contradictory statements.
+Furthermore, out-of-context errors are crafted by pairing $\mathcal{E}_{i}$ with a mismatched caption $c^{+}_{j}$ from another chart, where $i\neq j$.
+This simulates captions filled with unrelated data.
 The above process is illustrated in Algorithm 1.
@@
 Meta-evaluation of Different Evaluation Metrics
-Chocolate Model Lvlm Llm Ft LLaVA-1.5-13B 0.002 0.057 0.214 Bard -0.014 0.105 0.291 GPT-4V 0.157 0.205 0.215 DePlot + GPT-4 0.129 0.117 0.109 ChartVE (Ours) 0.178 0.091 0.215
+Chocolate Model Lvlm Llm Ft SummaC -0.011 0.023 0.036 QAFactEval 0.064 0.045 0.054 LLaVA-1.5-13B 0.002 0.057 0.214 ChartLlama 0.010 0.057 0.141 ChartAssistant-S 0.015 0.057 0.036 Bard -0.014 0.105 0.291 GPT-4V 0.157 0.205 0.215 DePlot + GPT-4 0.129 0.117 0.109 ChartVE (Ours) 0.178 0.091 0.215
 Table 3: Kendall’s Tau correlation of different approaches on the Chocolate dataset.
 To evaluate the effectiveness of different methods in assessing the factuality of generated captions on the Chocolate dataset, we employ Kendall’s Tau Kendall (1938) to compute the correlation between these methods and human judgments.
-Given the absence of prior work specifically focused on building factual inconsistency detection models for chart captions, we compare our ChartVE metric with methods that are capable of performing zero-shot reasoning including DePlot + GPT-4 Liu et al.
-(2023a); OpenAI (2023a), Bard Google (2023a), GPT-4V OpenAI (2023b), and the leading open-source LVLM, LLaVA-1.5-13B Liu et al.
-(2023c).
-The prompts for using these models as evaluation metrics can be found in Appendix E.
-The meta-evaluation results are summarized in Table 3.
-We observed that, overall, metrics exhibit the strongest correlation with human judgment on the Ft split and the weakest on the Lvlm split.
-This pattern aligns with expectations: the Ft captions are littered with more obvious mistakes, including a substantial number of out-of-context and nonsense errors, as detailed in Figure 1.
-Conversely, errors stemming from LVLMs often demand intricate inferences regarding the data points’ positions relative to the axes, thus presenting a more substantial challenge for detection.
-Importantly, our ChartVE metric demonstrates the highest correlation on the most demanding Lvlm split.
-However, ChartVE underperforms on the Llm split, potentially due to shifts in token distribution, as DePlot + GPT-4 occasionally employs table-centric terminology (e.g., “columns” and “entries”) absent from ChartVE’s training data.
-Despite this, ChartVE compares favorably to proprietary LVLMs and, notably, outstrips the performance of the SOTA open-source LVLM, LLaVA-1.5-13B, despite ChartVE being 100 times smaller in scale.
-On the Llm and Ft splits, the leading metrics are Bard and GPT-4V, respectively.
-Yet, what stands out is that Bard presents a negative correlation on the Lvlm split, signaling that LVLMs may falter significantly in evaluating the factuality of chart captions under certain conditions.
-As a result, we advocate for leveraging the best-performing metric for each split for evaluation.
+Given the absence of prior work on factual inconsistency detection methods for chart captions, we compare our ChartVE with zero-shot capable methods, including DePlot + GPT-4, Bard, GPT-4V, and the leading open-source LVLMs, LLaVA-1.5-13B Liu et al.
+(2023c), ChartLlama Han et al.
+(2023), and ChartAssistant-S Meng et al.
+(2024).
+Text-based factuality metrics, SummaC Laban et al.
+(2022) and QAFactEval Fabbri et al.
+(2022b), which compute the factual consistency between the reference caption and the generated caption, are also included.
+The prompts for these models are detailed in Appendix E.
+Meta-evaluation, summarized in Table 3, shows that, overall, metrics exhibit the strongest correlation with human judgment on the Ft split and the weakest on the Lvlm split.
+This pattern aligns with expectations: the Ft captions are littered with more obvious mistakes, such as out-of-context and nonsense errors, while errors stemming from LVLMs are harder to detect since they often demand intricate inferences regarding the data points’ positions relative to the axes, as detailed in Figure 1.
+Importantly, Our ChartVE excels on the challenging Lvlm split, but less so on the Llm split, likely due to shifts in token distribution, as DePlot + GPT-4 occasionally employs table-centric terminology (e.g., “columns” and “entries”) absent from ChartVE’s training data.
+Despite this, ChartVE compares favorably to proprietary LVLMs and outperforms open-source LVLMs, despite ChartVE being 64 times smaller in scale.
+Bard and GPT-4V lead on the Llm and Ft splits, respectively.
+However, Bard shows a negative correlation on the Lvlm split, hinting at LVLMs’ limitations in assessing the factuality of chart captions.
+Thus, we advocate for using the best-performing metric for each split for evaluation.
 4 Methodology
@@
 Our approach decomposes visual reasoning into image-to-structure rendering and text-based reasoning, allowing for interpretability and better correction of chart captions.
-To address the factual errors identified in generated captions, we propose a framework, C2TFec, that capitalizes a two-step correction process that offers interpretability.
-C2TFec first transforms input charts into structured data tables (Section 4.1) and then refines captions by rectifying any discrepancies using the tabular data (4.2).
-Our deep dive into the performance of a joint framework, “DePlot + GPT-4”, revealed that a notable proportion of errors in caption generation originated from the DePlot component.
-To mitigate this, we have developed a superior chart-to-table conversion model that builds upon a stronger base model, UniChart, vastly fortified with expansive fine-tuning datasets.
-The primary advantage of this refined methodology is its ability to harness the strong reasoning strengths of GPT-4 to faithfully correct errors, thereby significantly upscaling the factuality of the generated captions.
-[^6] Here, we do not consider approaches based on LVLMs due to their tendency towards factual errors as evidenced in Figure 1.
+In correcting factual errors in generated captions, we propose C2TFec, a two-step, interpretable framework, as shown in Figure 2.
+C2TFec first transforms input charts into data tables (Section 4.1), then rectifies errors in the caption using the tabular data (4.2).
+This framework is motivated by our analysis on “DePlot + GPT-4”, which shows that a notable proportion of errors in caption generation originated from the DePlot component.
+To mitigate this, we develop a stronger chart-to-table model based on UniChart, significantly improved with expansive fine-tuning datasets.
+The advantage of C2TFec is its ability to harness the reasoning strengths of GPT-4 to faithfully correct errors, boosting caption factuality.
+[^6] Here, we do not consider approaches based on LVLMs due to their tendency towards factual errors.
 […]
@@
 Levenshtein computes the edit distance between the corrected caption and the original caption (denoted as “N/A”).
-Metrics scores are shown separately for each of the three data splits based on captioning model source.
+Metric scores are shown separately for each of the three data splits based on captioning model source.
 The highest and second highest performing models per evaluation metric and split are highlighted in boldface and underlines respectively.
 4.1 Chart-To-Table Conversion
-The training data for our chart-to-table model is sourced from datasets including VisText Tang et al.
-(2023a), Chart-to-text Kantharaj et al.
-(2022), ChartQA Masry et al.
-(2022), and PlotQA Methani et al.
-(2020b), where we repurpose original charts and underlying data tables for our model’s training.
+The training data for our chart-to-table model is sourced from datasets including VisText, Chart-to-Text, ChartQA, and PlotQA, where we repurpose original charts and underlying data tables for our model’s training.
 We collected a total of 65K instances with a train:dev:test split of 61:2:2.
-In alignment with DePlot Liu et al.
+Similar to DePlot Liu et al.
 (2023a), our model is also trained to generate chart titles, enhancing its ability to contextualize the data represented in table form.
@@
 4.2 Table-based Error Rectification
-With the input chart now converted into structured tabular data, the second phase targets identifying and fixing factual inconsistencies in the initial caption $\mathcal{C}$.
-Our approach exploits the reasoning capacity of large language models, using GPT-4 to logically analyze the alignment between $\mathcal{C}$ and the constructed table $\mathcal{A}$.
-Specifically, GPT-4 first provides an explanatory breakdown of detected factual errors in $\mathcal{C}$ based on the table contents.
-It then uses this explanation to produce an edited caption $\mathcal{\hat{C}}$ remedying the identified inconsistencies.
+With the input chart now converted into structured tabular data, the second phase uses the reasoning capacity of LLMs to address the factual inconsistency between $\mathcal{C}$ and the generated table $\mathcal{A}$.
+Here, we use GPT-4 as the LLM.
+GPT-4 first provides an explanatory breakdown of detected factual errors in $\mathcal{C}$ based on the table contents.
+It then uses this explanation to produce a corrected caption $\mathcal{\hat{C}}$.
 This transparent process enables users to validate the reasoning behind each correction.
-Overall, the proposed C2TFec framework aims to produce verified chart captions by systematically isolating the language generation component from the analytical factual verification, taking advantage of the complementary strengths of separate vision and language models tailored to their respective domains.
-We hypothesize that the intermediate symbolic representation can act as an effective bridge to enforce and validate factual consistency.
+C2TFec separates the factual verification from language generation, taking advantage of the complementary strengths of separate vision and language models tailored to their respective domains.
+The symbolic table representation acts as a bridge to enhance and validate factual consistency in chart captions.
 5 Experimental Settings
-To evaluate our proposed C2TFec for factual error correction in chart captions, we conduct experiments on our annotated Chocolate dataset.
+To assess C2TFecs ability in factual error correction for chart captions, we experiment on the Chocolate dataset.
 Datasets
-We utilize the 1,187 chart-caption pairs from the Chocolate dataset, comprising charts paired with both factually consistent and inconsistent captions generated by various models.
-The captions contain a diverse range of factual errors annotated by human judges, allowing comprehensive assessment.
-As described in Section 2, the dataset is divided into Lvlm, Llm, and Ft splits based on the caption models.
+Our Chocolate dataset includes 1,187 chart-caption pairs with factually consistent and inconsistent captions, as detailed in Section 2.
+It is split into Lvlm, Llm, and Ft, reflecting the diversity of models that generated the captions.
 Baselines
-Since Chocolate does not comprise of training data, we compare C2TFec against competitive baselines that are capable of performing strong zero-shot reasoning.
-This includes state-of-the-art open-source and proprietary LVLMs, LLaVA-1.5-13B Liu et al.
-(2023c), GPT-4V OpenAI (2023b), and Bard Google (2023a), as well as a LLM-based pipeline, DePlot + GPT-4 Liu et al.
-(2023a); OpenAI (2023a).
+Since Chocolate does not comprise training data, we compare C2TFec against zero-shot capable LVLMs and LLMs, including LVLMs, LLaVA-1.5-13B, GPT-4V, Bard, as well as DePlot + GPT-4.
 For a fairer comparison between our approach and DePlot, we continue fine-tuning DePlot for an additional 5,000 steps on VisText, an approach which has been shown effective for adapting models to unseen domains Huang et al.
@@
 Evaluation Metrics
-We assess the factual consistency between corrected captions and input charts using the ChartVE metric, as well as GPT-4V and Bard, according to our recommendations in Section 3.2.
-In addition, since we require corrections to be made with as few edits as possible, we measure the number of edits using the Levenshtein distance Levenshtein et al.
+We assess the factual consistency between corrected captions and input charts using ChartVE, GPT-4V, and Bard, according to our recommendations in Section 3.2.
+In addition, since corrections should be made with as few edits as possible, we measure the number of edits using the Levenshtein distance Levenshtein et al.
 (1966).
@@
 6.1 Main Results
-The main results presented in Table 4 demonstrate that our proposed C2TFec framework achieves the best performance in terms of factual consistency on two out of three splits of the Chocolate dataset.
-Specifically, Chocolate attains the highest scores on the Lvlm and Ft splits, and achieves the second-best performance on the LLM.
+The results in Table 4 demonstrate that our C2TFec achieves the best performance for factual consistency on the Lvlm and Ft splits, and takes the second place on the Llm split.
 This indicates that the two-step process of first transforming charts into structured data tables and then rectifying factual inconsistencies using table-caption alignment is an effective strategy.
-An interesting observation is that Bard struggles on the Lvlm split.
-Combined with the negative correlation with human judgments of factuality that we observed in Table 3, we can conclude that Bard is unreliable in identifying factual errors in chart captions.
-Additionally, when used as an evaluator, GPT-4V tends to assign unduly positive factuality scores to its own corrected outputs.
-On the Lvlm and Ft splits, GPT-4V rates its own performance as significantly higher than all other models (see Table 6), while other metrics show GPT-4V lagging behind C2TFec.
+Additionally, we see that C2TFec outperforms the pipeline approaches of DePlot/$\text{DePlot}_{\text{CFT}}$ + GPT-4 across the board.
+While both methods utilize an intermediate tabular representation and leverage GPT-4 for language generation/correction, C2TFec employs a superior chart-to-table conversion model with much more comprehensive training datasets.
+This results in extracted tables that more faithfully capture the underlying chart data, better facilitating the downstream factual error correction.
+C2TFec also requires a relatively small number of edits to captions according to Levenshtein distance, making focused changes to improve factuality while minimizing revisions.
+An example output from C2TFec is shown in Figure 4.
+By comparison, the proprietary LVLM Bard produces corrected captions requiring 127.83 as many character-level edits on average.
+This signals excessive rewriting rather than targeted error correction.
+After manually inspecting Bard’s outputs, we found the reason is that Bard oftentimes try to improve the fluency of the caption by paraphrasing.
+Hence, it makes more edits to the generated captions.
+Figure 3: Human evaluation results on subsets of the Chocolate dataset, comparing C2TFec and GPT-4V.
+C2TFec corrects significantly more errors compared to GPT-4V, especially Value, Label, and Trend Errors.
+Bard’s underperformance on the Lvlm split and its negative correlation with human judgments of factuality, as shown in Table 3, implies its unreliability in detecting errors in chart captions.
+Additionally, when used as an evaluator, GPT-4V tends to assign high factuality scores to its own corrected outputs on all three splits (see Table 12), while other metrics show GPT-4V lagging behind C2TFec.
 This suggests GPT-4V may suffer from the self-enhancement bias Zheng et al.
-(2023), failing to reliably identify factual errors in its own generations when served as an evaluation metric.
-Hence, to further verify the effectiveness of our proposed approach, we conduct human evaluations in Section 6.2.
+(2023), overestimating its own performance when used for evaluation.
+We thus perform human evaluations in Section 6.2 to verify the effectiveness of our approach.
 6.2 Human Evaluation
-Our human evaluations aim to compare the performance of C2TFec and GPT-4V.
-The annotation tasks for human evaluation are the same as those presented to the annotators in Section 2: identifying various types of factual errors.
-Hence, the evaluation are conducted by the same annotators that perform the error analysis in Section 2.
+Our human assessments focus on comparing C2TFec with GPT-4V by using the same annotation tasks detailed in Section 2 for factual error identification, with the same annotators evaluating.
 We sampled 30 charts from each split of Lvlm, Llm, and Ft.
 For each chart, human judges are presented with a caption generated by one of the models.
-Figure 3 encapsulates the conclusive edge of C2TFec over GPT-4V across every error category.
-Notably, C2TFec demonstrates a marked reduction in Value Errors, surpassing GPT-4V by over a 20% margin within the Lvlm and Llm splits, and also half the overall error rate.
-Equally impressive was the near elimination of Trend Errors post-correction with C2TFec This underscores C2TFecs robust capability for accurate error rectification pertaining to reasoning regarding axes, including Label, Value, and Trend errors.
-A qualitative example is shown in Figure 4 to provide intuition on the superior correction ability of C2TFec over GPT-4V.
-We have also observed that the reason GPT-4V cannot correct these types of errors well is due to its inability to deduce the values of the data points from the input chart (see Figure 7).
-In contrast, GPT-4V exhibited a comparative proficiency in addressing Out-of-context Errors, involving information extraneous to the chart’s scope.
+Figure 3 demonstrates C2TFecs superiority in multiple error categories, especially with a substantial decrease in Value Errors, over 20% better in the Lvlm and Llm splits, and halving the overall error rate compared to GPT-4V.
+C2TFec virtually eliminated Trend Errors, highlighting its strong error correction ability, particularly for axes-related errors like Label, Value, and Trend errors.
+A representative comparison is shown in Figure 4.
+GPT-4V’s shortcomings seem to stem from its failure to accurately infer data point values from charts as evidenced in Figure 7.
+In contrast, GPT-4V is better in addressing Out-of-context Errors, involving information out of the chart’s scope.
 However, GPT-4V seemed challenged in rectifying errors within captions generated by itself, particularly within the Lvlm split.
@@
 Consequently, although GPT-4V’s capabilities are formidable, we recommend not using them to assess their own outputs.
-Figure 3: Human evaluation results on subsets of the Chocolate dataset, comparing C2TFec and GPT-4V.
-C2TFec corrects significantly more errors compared to GPT-4V, especially Value, Label, and Trend Errors.
 Figure 4: An example showing how decomposing the visual reasoning process into image-to-structure rendering and text-based reasoning allows C2TFec to accurately rectify errors in chart captions.
@@
 (2018), PlotQA Methani et al.
-(2020c), VisText Tang et al.
+(2020), VisText Tang et al.
 (2023a), and Chart-to-Text Kantharaj et al.
@@
 7.2 Factual Error Correction
-Previous research on based factual error correction has focused on two textual tasks: text summarization and fact-checking.
-Within text summarization, the bulk of work, such as Dong et al.
-(2020), has been carried out in the news domain and often involves methods that substitute inconsistent entities from the source text.
+Prior research in factual error correction has mainly targeted text summarization and fact-checking.
+Within summarization, the bulk of work has been carried out in the news domain and often involves methods that substitute inconsistent entities from the source text.
 Some studies have enhanced this approach through entity-replacement reranking techniques Chen et al.
@@
 In contrast, Fabbri et al.
-(2022) employed sentence compression datasets to train their models.
+(2022a) employed sentence compression datasets to train their models.
 More recently, Gao et al.
@@
 (2023d, c); Qiu et al.
+(2023b); Huang et al.
+(2024); Qiu et al.
 (2024).
@@
 Recently, Huang et al.
-(2023b) propose an interpretable framework that breaks down the process of fact-checking into individual components.
+(2023b) propose an interpretable framework that breaks down the process of factual error correction into individual components.
 Our study builds on these insights and extends them to a multimodal context, which challenges models to understand the chart images and the consistency between different modalities.
 8 Conclusion
-Our study exposes the prevalent issue of factual errors in chart captions generated by LVLMs, such as GPT-4V, and introduces Chocolate to scrutinize these errors.
+Our study exposes the prevalent issue of factual errors in chart captions generated by various chart captioning models and introduces Chocolate to scrutinize these errors.
 We establish the Chart Caption Factual Error Correction task to propel the creation of trustworthy captioning systems and present ChartVE, an evaluation model surpassing LVLMs in mirroring human assessments of caption factuality.
@@
 Future directions include extending our approach to multimodal contexts beyond charts, developing more sophisticated error detection and correction algorithms, and creating datasets covering a broader range of visual content.
+9 Ethical Considerations
+Text generation models pre-trained on information from the Web are known to demonstrate various biases.
+Despite the primary focus on models and datasets that represent the English-speaking population’s culture, manual examinations of the Chocolate dataset reveal no evidence of biases related to gender, age, race, or other socioeconomic factors Qiu et al.
+(2023a); Wang et al.
+(2024).
+In Section 2 and Section 6.2, we recruited annotators to assess the factual consistency of chart captions.
+The annotators were fairly compensated for their efforts, as detailed in Appendix B.
+During the annotation process, we made provisions for open communication, allowing the annotators the flexibility to work at their preferred pace and the freedom to withdraw from the project at any point.
+Additionally, we took measures to protect the anonymity of the contributors by excluding any personally identifiable information from the dataset.
 10 Limitations
@@
 Due to the scope of our study, we leave the prompt sensitivity experiments for future work.
-In addition, most charts in the datasets we used are line plots and bar plots.
-Future efforts can extend our work with additional analyses for other types of charts, such as scatter plots, violin plots, and distribution plots.
+In addition, charts in the datasets we used are mostly line plots and bar plots.
+Future efforts can extend our work with additional analyses for other types of charts, such as violin plots and distribution plots.
+Moreover, our investigation centered on the factuality of machine-generated chart captions, particularly those produced by LVLMs.
+We chose not to examine captions written by humans, as our primary objective was to highlight concerns regarding the reliability of automated chart captioning, a tool on which there is an increasing dependence for humans.
+The analysis of human-generated chart captions represents another avenue for future research, which could offer valuable comparisons and insights into the effectiveness of automated versus human captioning practices.
+Acknowledgement
+This research is based upon work supported by U.S.
+DARPA SemaFor Program No.
+HR001120C0123, DARPA ECOLE Program No.
+HR00112390060, and the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No.
+2019897 and No.
+2034562.
+The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
+Hou Pong Chan was supported in part by the Science and Technology Development Fund, Macau SAR (Grant Nos.
+FDCT/060/2022/AFJ, FDCT/0070/2022/AMJ) and the Multi-year Research Grant from the University of Macau (Grant No.
+MYRG2020-00054-FST).
 Appendix A Further Discussions
+Error Typology
+Our error typology was derived from our systematic preliminary analysis of around 50 generated captions.
+This exploratory phase allowed us to identify common error patterns and categories emergent from the data itself.
+Moreover, our methodology was enriched by drawing on insights from prior works such as Chart-to-Text, ChartT5, and VisText.
+Model Usage
+The GPT-4V version we used was gpt-4-vision-preview.
+We query GPT-4V via API and the cost for using GPT-4V to produce captions for our CHOCOLATE dataset is less than \$50 USD.
+For the Bard model, we obtain its outputs during October 2023.
+There is no cost using Bard since we access it via its web interface.
+Both models are prompted in single conversation, as seen in Figure 10.
 Captioning Model Analysis
@@
 Both observations could be explained by the fact that these models are sometimes confused about the charts in Pew, which are often associated with more complicated structures.
-Main Results
-We see that C2TFec outperforms the pipeline approaches of DePlot/$\text{DePlot}_{\text{CFT}}$ + GPT-4 across the board.
-While both methods utilize an intermediate tabular representation and leverage GPT-4 for language generation/correction, C2TFec employs a superior chart-to-table conversion model with much more comprehensive training datasets.
-This results in extracted tables that more faithfully capture the underlying chart data, better facilitating the downstream factual error correction.
-C2TFec also requires a relatively small number of edits to captions according to Levenshtein distance, making focused changes to improve factuality while minimizing revisions.
-An example output from C2TFec is shown in Figure 4.
-By comparison, the proprietary LVLM Bard produces corrected captions requiring 127.83 as many character-level edits on average.
-This signals excessive rewriting rather than targeted error correction.
+[…]
+Table 5: Sentence-level error counts and error rates.
+[…]
+Table 6: Caption-level error counts and error rates.
+Furthermore, in Figure 1, the error rates are computed as the number of such errors divided by the number of sentences.
+While this provides an overview of the frequency for each error, it does not indicate the likelihood of a value/label/trend/magnitude-related mention in the generated captions being factual.
+This limitation can result in an underrepresentation of certain error types – for instance, the infrequent occurrence of Magnitude Errors as shown in Figure 1 is more a consequence of the scarcity of magnitude-related mentions in the captions rather than an indication of the models’ superior trend variance comprehension.
+To address this, we sample 30 generated captions for each model from each dataset and compute another error rate as the number of sentences containing such non-factual mentions over the number of sentences containing such mentions.
+The results are shown in Table 7.
+The outcomes corroborate the observations in Section 2.2, while Table 7 offers a supplementary perspective on model performance.
+Meta-evaluation Results
+For the text-based metrics presented in Table 3, they both perform weakly in determining the factuality of the generated caption.
+This is largely because charts often contain much denser information compared to the corresponding reference.
+As a result, text-only factuality metrics are unsuitable for assessing factual consistency between charts and captions.
+Additionally, the negative values in Table 3 signal instances where the performance of the methods inversely correlates with human judgments, indicating their poor effectiveness in serving as evaluation metrics for identifying factual errors in generated captions.
+Furthermore, we experimented with ROC AUC as an additional meta-evaluation metric.
+As shown in Table 9, our original conclusion with Kendall’s Tau still holds.
+Understanding The Upper Bound
+We seek to understand the performance upper bound of our proposed two-stage framework by replacing generated tables with ground-truth data tables.
+Since the ground-truth data tables in Pew are not available, we experiment with only the instances from the VisText dataset.
+The results are demonstrated in Table 8.
+Ablation Studies
+We compare our chart-to-table component against DePlot, $\text{DePlot}_{\text{CFT}}$, and GPT-4V.
+Since the Pew split of Chart-to-Text does not contain ground truth labels, we show performance on the VisText test set.
+We use RMS-F1 proposed in Liu et al.
+(2023a) as the evaluation metric.
+From Table 11, we see that our Chart-to-Table component significantly outperforms all baselines, indicating its effectiveness in converting charts into their underlying data tables.
+Figure 5: Human annotation interface for our data collection discussed in Section 2.
+Examples of each type of error from Table 1 are also displayed in the annotation interface.
+We were not able to show these examples in this figure due to space limits.
 Appendix B Annotation Details
@@
 Additionally, each HIT was meticulously crafted to ensure that annotators could achieve an equivalent hourly pay rate of \$15 - \$20, assuming they work without interruption.
+[…]
+Table 7: Error rates (%) are calculated by dividing the number of sentences containing such non-factual mentions (e.g.
+non-factual mentions of values) by the number of sentences containing such mentions (e.g.
+all mentions of values).
+The lower the error rate, the better the performance.
+[…]
+Table 8: Correction performance of different models on the Chocolate dataset.
+ChartVE measures factuality by computing the entailment probability from each chart to the corresponding caption sentences.
+GPT-4V and Bard, when used as evaluation metrics, rate each chart caption as factually consistent with the chart or not.
+Levenshtein computes the edit distance between the corrected caption and the original caption.
+Metric scores are shown separately for each of the three data splits.
+Note that the Bard metric corresponds to Gemini Pro Google (2023b) since the experiments were conducted after its release.
+Chocolate Model Lvlm Llm Ft LLaVA-1.5-13B 0.501 0.566 0.670 Bard 0.488 0.617 0.743 GPT-4V 0.638 0.738 0.678 DePlot + GPT-4 0.609 0.629 0.590 ChartVE (Ours) 0.646 0.595 0.676
+Table 9: ROC AUC of different approaches on the Chocolate dataset.
 B.2 Annotation Guidelines
@@
 Chocolate-Lvlm Chocolate-Llm Chocolate-Ft # Factual # Non-factual # Factual # Non-factual # Factual # Non-factual Sentence 1,683 1,270 518 469 360 1,023 Caption 74 321 27 169 112 484
-Table 5: Dataset statistics per pplit. A sentence is considered factual if and only if it does not contain any factual error. A caption is considered factual if all its sentences are factual.
+Table 10: Dataset statistics per split. A sentence is considered factual if and only if it does not contain any factual error. A caption is considered factual if all its sentences are factual.
 Appendix D Implementation Details
+1
 Input: Data table $\mathcal{A}_{\mathcal{E}_{i}}$ for chart $\mathcal{E}_{i}$, Positive caption sentence $c^{+}_{i}$.
 Output: Set of negative caption sentences $C^{-}_{i}=\{c^{-}_{i,\text{value}},c^{-}_{i,\text{trend}},c^{-}_{i,\text{context}}\}$.
-1 Initialize $C^{-}_{i}$ as an empty set;
+2
+3Initialize $C^{-}_{i}$ as an empty set;
 4 Define a lexicon of trend terms $T$;
@@
 11             Add $c^{-}_{i,\text{value}}$ to $C^{-}_{i}$;
-10
-11
-12// Generate Trend Errors;
+12
+13
+14// Generate Trend Errors;
 15 for each trend term $t$ in $T$ do
@@
 18             Add $c^{-}_{i,\text{trend}}$ to $C^{-}_{i}$;
-17
-18
-19// Generate Out-of-Context Errors;
+19
+20
+21// Generate Out-of-Context Errors;
 22 Randomly select a different chart $\mathcal{E}_{j}$ where $j\neq i$;
@@
 24 Add $c^{-}_{i,\text{context}}$ to $C^{-}_{i}$;
-23 return $C^{-}_{i}$;
+25
+26return $C^{-}_{i}$;
+27
 Algorithm 1 Table-guided Negative Data Generation
+Model RMS-F1 DePlot 15.97 $\text{DePlot}_{\text{CFT}}$ 78.23 GPT-4V 10.44 Chart-to-Table (Ours) 83.60
+Table 11: Chart-to-Table performance comparison on the VisText dataset.
 D.1 Details of the Chart-To-Table Model
@@
 Figure 7: An example showing GPT-4V cannot accurately extract tables from charts. This indicates its inability to infer the actual value of each data point within the chart.
-Appendix E Prompts Used In Experiments
+Appendix E Prompts
 The prompts for using LVLM and LLM as evaluation metrics are displayed in Figure 8 and Figure 9, while the prompts for factual error correction are shown in Figure 10 and Figure 11.